# Incommensurate Twisted Bilayer Graphene: emerging quasi-periodicity and stability

### v1

@@ III Perturbative expansion and small divisors
To compute the Schwinger functions $S(\mathbf{k})$, we will use Feynman graph expansions. We give the rule setting $\nu=0$ for definiteness. The graphs for this model are chain graphs of the form depicted in Figure 1
@@
Each line $s$ is associated a layer label $i_{s}\in\{1,2\}$ and two valley indices $\alpha_{s},\alpha^{\prime}_{s}\in\{a,b\}$. Each vertex has one entering line $s_{1}$ and an exiting line $s_{2}$, and we impose that these lines have different indices: $i_{s_{1}}=3-i_{s_{2}}$, and the same valley indices $\alpha_{s_{1}}=\alpha_{s_{2}}$. To each vertex $r$ that has an entering line $s_{1}$ and exiting line $s_{2}$, we associate an index, which if $i_{s_{1}}=1$ we denote by $l_{s}\in\mathbb{Z}^{2}$, and if $i_{s_{2}}=2$ we denote by $m_{s}\in\mathbb{Z}^{2}$.
@@
The value of a graph is obtained by taking the product over the lines of the corresponding propagators, and multiplying them by the product over the vertices of the interaction terms. The Schwinger function is then obtained by taking a sum over all possible graphs with the correct external labels. A more explicit expression for the Schwinger function is given in Appendix C.
@@
The persistence or not of the semimetallic behaviour depends on the convergence of the expansion. One of the more important and difficult reasons for which convergence could, in principle, not occur, is that small divisors could accumulate, as we will now explain. Note that if ${\bf k}_{1}$ and ${\bf k}_{2}$ are 2 neighboring terms their difference is given by $k_{1}-k_{2}+mb^{\prime}+lb$; moreover the propagator are singular at the Dirac points $p_{F}^{\omega,i}$. Consider therefore $g_{1}({\bf k})\tau^{(1)}_{l}(k+lb)g_{2}({\bf k}+lb)$ and suppose that $k$ is in the first Brillouin zone and is close to $p_{F,1}^{\omega}$, that is $k=p_{F,1}^{\omega}+r_{1}$ with $r_{1}=O(\varepsilon)$ and $\varepsilon$ is a small parameter (so that the propagator is $O(\varepsilon^{-1})$); suppose now that $k+lb+mb^{\prime}$ is also in the first Brillouin zone (by periodicity we can always add $mb^{\prime}$ to achieve this) and close to $p_{F,2}^{\omega^{\prime}}$, that is $k+lb+mb^{\prime}=p_{F,2}^{\omega}+r_{2}$ with $r_{2}=O(\varepsilon)$. In principle this would produce an $O(\varepsilon^{-2})$ contribution, that is an accumulation of small divisors which could produce large contributions which can destroy convergence. Note however that
$O(\varepsilon)=|r_{1}|+|r_{2}|\geq|r_{1}-r_{2}|=|lb+mb^{\prime}-p_{F,2}^{\omega^{\prime}}-p_{F,1}^{\omega}|$ (27)
@@
so that this accumulation of small divisors is possible only when the quantity in the r.h.s. is small. The effect of such terms is rather delicate to be understoord. Indeed similar terms in the case of random potential produce a localization phase in 1d, while an extended phase in hgher dimension, while a quasi periodic disorder in 1d produces an extended phase.
IV Diophantine condtions
@@
As we noted above, the accumulation of small divisors can only occur when $|lb+mb^{\prime}+p_{F,i}^{\omega}-p_{F,j}^{\omega^{\prime}}|$ is small. For generic values of $\theta$ (for which $s_{\theta}$ or $c_{\theta}$ is irrational), this happens when $l,m$ are large enough, and the basic issue for stability is if small divisors are balanced by the fact that the interlayer hopping is weak for large values of $l,m$. We therefore need to quantify the relation between the size of $lb+mb^{\prime}+p_{F,i}^{\omega}-p_{F,j}^{\omega^{\prime}}$ and that of $l,m$, which we do using number theoretical considerations. One such consideration is the so called Diophantine condition, which consists in restricting ourselves to values of $\theta$ for which, for any $\omega,\omega^{\prime},i,j$
$|p_{F,i}^{\omega}-p_{F,j}^{\omega^{\prime}}+lb+mb^{\prime}|\geq{C_{0}\over|y|^{\tau}}.$ (28)
@@
where $y$ is either $y=l$ or $y=m$ and $y\neq 0$, If this is satisfied, then when $|p_{F,i}^{\omega}-p_{F,j}^{\omega^{\prime}}+lb+mb^{\prime}|=O(\epsilon)$, then
@@
Therefore $|l|\geq\varepsilon^{-1/\tau}$, and, due to the exponential decay of $\widehat{\varsigma}$ in $\tau^{(1)}$ (see (12)), so $|\tau^{(1)}_{l}(k+lb)|\leq e^{-\kappa\varepsilon^{-1/\tau}}$ which compensates the small divisors $\varepsilon^{-2}$. This simple argument says that small divisors due to adjacent propagators do not accumulate. Of course this argument is not sufficient to prove the convergence of the series, and then the persistence of cones; it says only that two adjacent propagators cannot be simultaneously small but it says nothing about non adjacent ones, which is the generic case. This requires a Renormalization Group analysis, as we will discuss below.
A basic preliminary question is whether there are $\theta$’s such that (28) holds. Such condition is usually assumed in KAM theory or in the case of fermions with quasi periodic potentials; in such cases the frequencies are independent so that the fact that there is a large measure set of them verifying (28) follows from standard notions of number theory. In the present case however the issue is much more subtle as all the frequencies depend on a single parameters $\theta$. Nevertheless, we prove the following lemma.
Lemma IV.1.
For every interval $\theta_{0},\theta_{1}]\subset[0,2\pi)$, the set ${\cal D}:=\{\theta\in[\theta_{0},\theta_{1}]$ satisfying (28)}\} has measure $1-O(C_{0}/(\theta_{1}-\theta_{0})^{2})$, so that choosing $C_{0}$ small enough with respect to $\theta_{1}-\theta_{0}$, $\mathcal{D}$ has large relative measure.
We will provide a full proof of this lemma in Appendix B. For the moment, let us discuss the main idea of this proof. In order to keep things simple, we will focus on the case $i=j$ and $\omega=\omega^{\prime}$ here. Let $M:=lb+mb^{\prime}$ and we wish to obtain a lower bound on $|M|$. To do so, we use a simple inequality: $\forall x,y\in\mathbb{R}$,
@@
$\sum_{m,l_{\omega}}^{*}\int_{-C_{1}|m|^{-\tau}}^{C_{1}|m|^{-\tau}}\frac{1}{g_{l_{\omega},m}^{\prime}}dg_{l_{\omega},m}$ (34)
@@
$\sum^{*}_{m}2C_{1}{|m|^{1-\tau}\over\min_{\theta}|m\cdot f_{\omega}^{\prime}(\theta)|}$ (35)
@@
Now, in order for this bound to be useful, we need a good bound on $m\cdot f_{\omega}^{\prime}$. Now $m\cdot f_{\omega}^{\prime}(\theta)$ can be small, but only if $m$ is large enough. This suggests we should control it using another Diophantine condition, but we would run into an infinite problem: we would need $mf_{\omega}^{\prime\prime}$ to be bounded below, for which we would impose an extra Diophantine condition, which would itself require a Diophantine condition on the third derivatives, et cætera. We can avoid this problem as follows. If $f_{\omega}^{\prime}(0)=|f_{\omega}^{\prime}(0)|(\cos\beta,\sin\beta)$ is non vanishing and $\theta$ is small then $(m\cdot f_{\omega}^{\prime}/|f_{\omega}^{\prime}|)\sim|m|\cos(\theta_{m})$ where $\theta_{m}$ is the angle between $\beta$ and $m$. We can distinguish in the sum (35) the sum over $m$ such that $|\theta_{m}|,|\theta_{m}-\pi|<\pi/4$ and the complementary set $|\theta_{m}-\pi/2|,|\theta_{m}-3\pi/2|<\pi/4$. In the first sum, $(m\cdot f_{\omega}^{\prime}/|f_{\omega}^{\prime}|)$ will be greater than $|m|$ up to some constant; we impose a Diophantine condition only for the second term: for $|\theta_{m}-\pi/2|,|\theta_{m}-3\pi/2|<\pi/4$ we assume that $|l_{\omega}+m\cdot f_{\omega}^{\prime}|\leq C_{1}|m|^{-\tau}$. Again we will end-up with a condition like (35) involving $m\cdot(f_{\omega}^{\prime}/|f_{\omega}^{\prime}|)^{\prime}$ which in principle could be arbitrarily small. However $(f^{\prime}/|f^{\prime}|)^{\prime}$ is orthogonal to $f_{\omega}^{\prime}/|f_{\omega}^{\prime}|$ hence $m\cdot(f_{\omega}^{\prime}/|f_{\omega}^{\prime}|)^{\prime}\sim|m|\cos(\theta_{m}+\pi/2)$ which, in this region, is greater than $|m|$.
@@
Thus, we construct a large-measure set of $\theta$’s that satisfy $|l_{\omega}+m\cdot f_{\omega}(\theta)|\geqslant C_{1}|m|^{-\tau}$. A detailed proof can be found in Appendix B.
V Stability of the semimetallic phase
We are now ready to state our main result.
Theorem For any interval $\theta_{0},\theta_{1}]\subset[0,2\pi)$, any $C_{0}>0$, there exists a subset of $\theta_{0},\theta_{1}]$ whose measure is at least $1-O(C_{0}/(\theta_{0}-\theta_{1})^{2})$ such that (28) holds, and an $\varepsilon_{0}$ (depending on $C_{0},\theta_{0},\theta_{1}$), such that, for any $|\lambda|\leq\varepsilon_{0}$, for a suitable choice of $\nu_{i,\omega}$ $(\widehat{S}_{j,j}(\mathbf{k}+\mathbf{p}_{F,j}^{\omega}))=$
$\begin{pmatrix}-iZ_{j,\omega}k_{0}&(iv_{j,\omega}k_{1}-w_{j,\omega}\omega k_{2})\\ (-iv_{j,\omega}^{*}k_{1}-w_{j,\omega}^{*}\omega k_{2})&-iZ_{j,\omega}k_{0}\end{pmatrix}^{-1}(1+O(|k|^{\alpha}))$ (36)
with $0\leq\alpha\leq 1$, $Z=1+O(\lambda)$ real and $v_{i,\omega}=3t/2+O(\lambda)$, $w_{i,\omega}=3t/2+O(\lambda)$, $\nu_{j,\omega}=O(\lambda)$.
This result ensures that, even taking into account the Umklapp processes involving the exchange of very high momenta due to the emerging quasi-periodicity, the Weyl semimetallic phase persists for small interalyer coupling and a large measure set of angles.
The interlayer coupling modifies the position of the Dirac points; we have properly chosen the bare Dirac points $p_{F,i}^{\pm}(\lambda)$ in absence of interlayer coupling given by $\Omega(p_{F,i}^{\pm}(\lambda))+\nu_{i,\pm}=0$) so that their renormalized physical value is $p_{F,i}^{\pm}$ given by (10). This is essentially equivalent to say that the position of the Dirac points genericaly moves in a way depending on the angle, the layer and the coupling.
The velocities $w_{j,\omega},v_{j,\omega}$ and the wave function normalization $Z_{j,\omega}$, are renormalizated in a way generically dependent on the layer and the angle. Note that a priori several other relevant terms could be present, but they are excluded by symmetry. The singularity of the Schiwnger function is given by $Z^{2}k_{0}^{2}+R(k)$ with $R(k)\sim(|v|^{2}k_{1}^{2}+|w|^{2}k_{2}^{2})$; the singularity of the 2-point function is therefore the same as in absence of interlayer at weak coupling ensuring the stability of the semimetallic phase.
The result holds for irrational twisting angles verifying (28). The relative measure of this set can be made arbitrarely close to $1$ by decreasing $C_{0}$. In the remaining sections we prove the above result by a Renormalization Group analysis, leading to a convergent expansion.
@@ VI.1 Multiscale decomposition
We introduce smooth cut-off functions: for $i=1,2$, $\omega=\pm$, $h\in\{-\infty,\cdots,0\}$, let $\chi_{h,i,\omega}({\bf k})$ be a smooth function that vanishes outside the region ${\bf k}-{\bf p}_{F,i}^{\omega}||\leq\gamma^{h-1}$ and that is equal to 1 for $k-{\bf p}_{F,i}^{\omega}||\geq\gamma^{h-2}$. The constant $\gamma>1$ will be chosen below to be large enough. Note that, in this way, the supports of $\chi_{0,i,+}$ and $\chi_{0,i,-}$ do not overlap. We define $\widehat{g}_{i,\omega}^{(\leq 0)}({\bf k})=\chi_{0,i,\omega}({\bf k})\widehat{g}_{i}({\bf k})$ and
@@
with $\widehat{g}^{(1)}({\bf k})=(1-\sum_{\omega}\chi_{0,i,\omega}(k))\widehat{g}_{i}({\bf k})$; this induces the Grassmann variable decomposition $\widehat{\psi}_{i,{\bf k},\alpha}=\widehat{\psi}_{i,{\bf k},\alpha}^{(1)}+\sum_{\omega=\pm}\widehat{\psi}^{(\leq 0)}_{i,{\bf k},\alpha,\omega}$ with propagators given by $\widehat{g}^{(1)}_{i}$ and $\widehat{g}^{(\leq 0)}_{i,\omega}$ respectively. Note that $\widehat{\psi}^{(1)}$ correspond to fermions with momenta far from the Fermi points, while $\widehat{\psi}^{(\leq 0)}$ with momenta around $\pm{\bf p}_{F,i}$.
@@
where $\widehat{g}_{i,\omega}^{(h)}({\bf k}):=f_{h,i,\omega}({\bf k})\widehat{g}_{i,\omega}^{(\leq 0)}$ in which $f_{h,i,\omega}:=\chi_{h,i,\omega}-\chi_{h-1,i,\omega}$ is a smooth cutoff function supported in $\gamma^{h-3}\leq|{\bf k}-{\bf p}^{\omega}_{F,i}|\leq\gamma^{h-1}$ such that $\sum_{h=-\infty}^{0}f_{h,i,\omega}=\chi_{0,i,\omega}$. The integration is done recursively in the following way: assume that we have integrated the fields $\psi^{(1)},..,\psi^{(h-1)}$ obtaining
$e^{W}=\int\bar{P}(d\psi^{(\leq h)})e^{V^{(h)}(\psi,\phi)}$ (39)
@@
$\sum_{i,\omega,\omega^{\prime},m,\alpha,\alpha^{\prime}}\int_{\widehat{\Lambda}_{i}}d{\bf k}W^{(h,\omega,\omega^{\prime})}_{i,1,m,\alpha,\alpha^{\prime}}({\bf k})\psi_{i,{\bf k},\alpha,\omega}^{+}\psi_{1,{\bf k}+mb^{\prime},\alpha^{\prime},\omega^{\prime}}^{-}.$ (41)
@@
${\cal L}W^{(h,\omega,\omega^{\prime})}_{i,j,l}({\bf k})=\delta_{\omega,\omega^{\prime}}\delta_{i,j}\delta_{l,0}[W^{(h,\omega,\omega)}_{i,i,0}(0,p_{F,i}^{\omega})+$
$k_{0}\partial_{0}W^{(h,\omega,\omega)}_{i,i,0}(0,p_{F,i}^{\omega})+(k-p_{F,i}^{\omega})\partial W^{(h,\omega,\omega)}_{i,i,0}(0,p_{F,i}^{\omega}).$ (42)
@@
${\cal L}_{2}V^{(h)}:=\sum_{i,j,\omega,\alpha,\alpha^{\prime}}\int_{\widehat{\Lambda}_{i}}d{\bf k}z_{h,\omega,\alpha,\alpha^{\prime},i,j}(\mathbf{k}-\mathbf{p}_{F,i}^{\omega})_{j}\psi^{+}_{{\bf k},\omega,i,\alpha}\psi^{-}_{{\bf k},\omega,i,\alpha^{\prime}}$ (45)
@@
$\nu_{h,\omega,\alpha,\alpha^{\prime},i}:=\gamma^{-h}W_{i,i,0,\alpha,\alpha^{\prime}}^{h,\omega,\omega}(0,p_{F,i}^{\omega})$ (46)
$z_{h,\omega,\alpha,\alpha^{\prime},i,j}:=-\partial_{\mathbf{k}_{j}}W_{i,i,0,\alpha,\alpha^{\prime}}^{h,\omega,\omega}(0,p_{F,i}^{\omega}).$ (47)

### v2

@@ III Perturbative expansion and small divisors
To compute the 2-point function $\widehat{S}(\mathbf{k})$, we will use Feynman graph expansions. The graphs for this model are chain graphs of the form depicted in Figure 2
@@
Let us give the rule first for the graphs with $\nu=0$. Each line $s$ is associated a layer label $i_{s}\in\{1,2\}$ and two indices $\alpha_{s},\alpha^{\prime}_{s}\in\{a,b\}$. Each vertex has one entering line $s_{1}$ and an exiting line $s_{2}$, and we impose that these lines have different indices: $i_{s_{1}}=3-i_{s_{2}}$, and the same indices $\alpha_{s_{1}}=\alpha_{s_{2}}$. To each vertex $r$ that has an entering line $s_{1}$ and exiting line $s_{2}$, we associate an index, which if $i_{s_{1}}=1$ we denote by $l_{s}\in\mathbb{Z}^{2}$, and if $i_{s_{2}}=2$ we denote by $m_{s}\in\mathbb{Z}^{2}$.
@@
The value of a graph is obtained by taking the product over the lines of the corresponding propagators, and multiplying them by the product over the vertices of the interaction terms. The 2-point function is then obtained by taking a sum over all possible graphs with the correct external labels. A more explicit expression for the 2-point function is given in Appendix C. The graphs with $\nu\not=0$ are obtained by the above ones in the standard way adding vertices such that the momentum of the external lines is the same.
@@
The persistence or not of the semimetallic behavior depends on the convergence of the expansion. One of the more important and difficult reasons for which convergence could, in principle, not occur, is that small divisors could accumulate, as we will now explain. Note that if ${\bf k}_{1}$ and ${\bf k}_{2}$ are 2 neighboring terms their difference is given by $k_{1}-k_{2}+mb^{\prime}+lb$; moreover the propagator is singular at the Dirac points $K_{\omega}^{i}$. Consider therefore $g_{1}({\bf k})\tau^{(1)}_{l}(k+lb)g_{2}({\bf k}+lb)$ and suppose that $k$ is in the first Brillouin zone and is close to $K^{1}_{\omega}$, that is $k=K^{1}_{\omega}+r_{1}$ with $r_{1}=O(\varepsilon)$ and $\varepsilon$ is a small parameter (so that the propagator is $O(\varepsilon^{-1})$); suppose now that $k+lb+mb^{\prime}$ is also in the first Brillouin zone (by periodicity we can always add $mb^{\prime}$ to achieve this) and close to $K^{2}_{\omega^{\prime}}$, that is $k+lb+mb^{\prime}=K^{2}_{\omega}+r_{2}$ with $r_{2}=O(\varepsilon)$. In principle this would produce an $O(\varepsilon^{-2})$ contribution, that is an accumulation of small divisors which could produce large contributions which can destroy convergence. Note however that
$O(\varepsilon)=|r_{1}|+|r_{2}|\geq|r_{1}-r_{2}|=|lb+mb^{\prime}-K^{2}_{\omega^{\prime}}+K^{1}_{\omega}|$ (31)
@@
so that this accumulation of small divisors is possible only when the quantity in the r.h.s. is small. The effect of such terms is rather delicate to be understood. Indeed similar terms in the case of random potential produce a localization phase in 1d, while an extended phase in higher dimension, while a quasi periodic disorder in 1d produces an extended phase.
IV Main result
@@
As we noted above, the accumulation of small divisors can only occur when $|lb+mb^{\prime}+K^{i}_{\omega}-K^{j}_{\omega^{\prime}}|$ is small. For generic values of $\theta$ (for which $s_{\theta}$ or $c_{\theta}$ is irrational), this happens when $l,m$ are large enough, and the basic issue for stability is whether small divisors are balanced by the fact that the interlayer hopping is weak for large values of $l,m$. We therefore need to quantify the relation between the size of $lb+mb^{\prime}+K^{i}_{\omega}-K^{j}_{\omega^{\prime}}$ and that of $l,m$, which we do using number-theoretical considerations. One such consideration is imposing on the angles a Diophantine condition (4). We will prove in Lemma V.1 below that this condition holds for most angles, in a sense made precise below.
@@
If this is satisfied, then when $|K^{i}_{\omega}-K^{j}_{\omega^{\prime}}+lb+mb^{\prime}|=O(\epsilon)$, we have
@@
Therefore $|l|\geq\varepsilon^{-1/\tau}$, and, due to the exponential decay of $\widehat{\varsigma}$ in $\tau^{(1)}$ (see (15)), we get $|\tau^{(1)}_{l}(k+lb)|\leq e^{-\kappa\varepsilon^{-1/\tau}}$, which compensates the small divisors $\varepsilon^{-2}$. This simple argument shows that small divisors due to adjacent propagators do not accumulate. Of course, this argument is not sufficient to prove the convergence of the series, and thus the persistence of cones; it only says that two adjacent propagators cannot be simultaneously small, but it says nothing about non-adjacent ones, which is the generic case. This requires a Renormalization Group analysis, as we will discuss below. We first present our main result.
Theorem IV.1.
For any $\theta$ chosen to satisfy the Diophantine condition (4), there exists an $\varepsilon_{0}(C_{0})$ such that, for any $|\lambda|\leq\varepsilon_{0}$, there exists $\nu_{j,\omega}$ such that, in the limit $\beta\rightarrow\infty$ the 2-point function verifies
$(\widehat{S}_{j,j}(\mathbf{k}+\mathbf{K}^{j}_{\omega}))=$
$\begin{pmatrix}-iZ_{j,\omega}k_{0}&(iv_{j,\omega}k_{1}-w_{j,\omega}\omega k_{2})\\ (-iv_{j,\omega}^{*}k_{1}-w_{j,\omega}^{*}\omega k_{2})&-iZ_{j,\omega}k_{0}\end{pmatrix}^{-1}(1+O(|\mathbf{k}|^{\alpha}))$ (33)
with $0\leq\alpha\leq 1$, $Z_{j,\omega}=1+O(\lambda)$ real and $v_{j,\omega}=3t/2+O(\lambda)$, $w_{j,\omega}=3t/2+O(\lambda)$, $\nu_{j,\omega}=O(\lambda)$.
The interlayer coupling modifies the position of the Dirac points; we have properly chosen the bare Dirac points $K^{j}_{\pm}(\lambda)$ in the absence of interlayer coupling (given by $\Omega(K^{j}_{\pm}(\lambda))+\nu_{j,\pm}=0$) so that their renormalized physical value is $K^{i}_{\pm}$ given by (13). This is equivalent to say that the position of the Dirac points generically moves depending on the angle, the layer, and the coupling.
The velocities $w_{j,\omega},v_{j,\omega}$ and the wave-function normalization $Z_{j,\omega}$ are renormalized in a way generically dependent on the layer and the angle. Note that a priori several other relevant terms could be present, but they are excluded by symmetry. The singularity of the 2-point function is given by $Z_{j,\omega}^{2}k_{0}^{2}+R(k)$ with $R(k)\sim(|v_{j,\omega}|^{2}k_{1}^{2}+|w_{j,\omega}|^{2}k_{2}^{2})$; the singularity of the 2-point function is therefore the same as in the absence of interlayer coupling at weak interaction, proving the stability of the semimetallic phase.
V The Diophantine condition
One has, of course, to check the validity of (4). As the value of the Diophantine constant $C_{0}$ is expected to depend strongly on the exact value of $\theta$, it is convenient to give a generic statement for angles in a certain set. As in the case of disordered systems, we give a statement in measure (or probability) sense. We want to estimate the relative measure of the set for which (4) holds. This requires some care as the $K_{\pm},b,b^{\prime}$ are not independent parameters but given functions of the angles.
As a starting point, let us consider the case $i=j$ and $\omega=\omega^{\prime}$. Let $M:=lb+mb^{\prime}$ and we wish to obtain a lower bound on $|M|$. To do so, we use a simple inequality: $\forall x,y\in\mathbb{R}$,
@@ V The Diophantine condition
$f_{\omega}(\theta):=\left(\frac{\varphi_{1}-\omega\varphi_{3}}{2},\frac{\varphi_{2}+\omega\varphi_{4}}{2}\right).$ (37)
@@
$\sum_{m,l_{\omega}}^{*}\int_{-C_{1}|m|^{-\tau}}^{C_{1}|m|^{-\tau}}\frac{1}{g^{\prime}_{l_{\omega},m}}\,dg_{l_{\omega},m}$ (38)
@@
$\sum^{*}_{m}2C_{1}\frac{|m|^{1-\tau}}{\min_{\theta}|m\cdot f_{\omega}^{\prime}(\theta)|}$ (39)
@@
Now, in order for this bound to be useful, we need a good bound on $m\cdot f_{\omega}^{\prime}$. Now $m\cdot f_{\omega}^{\prime}(\theta)$ can be small, but only if $m$ is large enough. This suggests we should control it using another Diophantine condition, but we would run into an infinite problem: we would need $mf_{\omega}^{\prime\prime}$ to be bounded below, for which we would impose an extra Diophantine condition, which would itself require a Diophantine condition on the third derivatives, et cætera.
We can avoid this problem as follows. Let us fix for definiteness $\theta$ in a small interval around $\bar{\theta}$. If $f_{\omega}^{\prime}(\bar{\theta})=|f_{\omega}^{\prime}(\bar{\theta})|(\cos\beta,\sin\beta)$ is non-vanishing and $\theta-\bar{\theta}$ is small, then $(m\cdot f_{\omega}^{\prime}/|f_{\omega}^{\prime}|)\sim|m|\cos(\theta_{m})$ where $\theta_{m}$ is the angle between $\beta$ and $m$. We can distinguish in the sum (39) the sum over $m$ such that $|\theta_{m}|,|\theta_{m}-\pi|<\pi/4$ and the complementary set $|\theta_{m}-\pi/2|,|\theta_{m}-3\pi/2|<\pi/4$. In the first sum, $(m\cdot f_{\omega}^{\prime}/|f_{\omega}^{\prime}|)$ will be greater than $|m|$ up to some constant; we impose a Diophantine condition only for the second term: for $|\theta_{m}-\pi/2|,|\theta_{m}-3\pi/2|<\pi/4$ we assume that $|m\cdot f_{\omega}^{\prime}|\leq C_{1}|m|^{-\tau}$. Again we will end up with a condition like (39) involving $m\cdot(f_{\omega}^{\prime}/|f_{\omega}^{\prime}|)^{\prime}$ which in principle could be arbitrarily small. However $(f^{\prime}/|f^{\prime}|)^{\prime}$ is orthogonal to $f_{\omega}^{\prime}/|f_{\omega}^{\prime}|$, hence $m\cdot(f_{\omega}^{\prime}/|f_{\omega}^{\prime}|)^{\prime}\sim|m|\cos(\theta_{m}+\pi/2)$ which, in this region, is greater than $|m|$. Note that the derivatives $f^{\prime}(\bar{\theta})$ must be finite.
@@
Thus, we construct a large-measure set of $\theta$’s that satisfy $|l_{\omega}+m\cdot f_{\omega}(\theta)|\geqslant C_{1}|m|^{-\tau}$. More exactly, in appendix B.2, we prove the following lemma (this proof is computer-assisted).
Lemma V.1.
For any $\theta\in[0,2\pi)\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2})$ where $\mathcal{B}_{1}\cup\mathcal{B}_{2}$ is a finite set of special angles (see (136) and (137) for an explicit expression), and any $\delta\theta$ small enough, for almost every angle in $\theta-\delta\theta,\theta+\delta\theta]$, there exists $C_{0}$ such that, for all $\omega,\omega^{\prime},i,j$, and for all $l,m\in\mathbb{Z}^{2}\setminus\{0\}$,
$|K^{i}_{\omega}-K^{j}_{\omega^{\prime}}+lb+mb^{\prime}|\geq\frac{C_{0}}{|m|^{\tau}},\quad|K^{i}_{\omega}-K^{j}_{\omega^{\prime}}+lb+mb^{\prime}|\geq\frac{C_{0}}{|l|^{\tau}}.$ (40)
@@ VI.1 Multiscale decomposition
We introduce smooth cut-off functions: for $i=1,2$, $\omega=\pm$, $h\in\{-\infty,\cdots,0\}$, let $\chi_{h,i,\omega}({\bf k})$ be a smooth function that vanishes outside the region ${\bf k}-\mathbf{K}^{i}_{\omega}||\leq\gamma^{h-1}$ and that is equal to 1 for ${\bf k}-\mathbf{K}^{i}_{\omega}||\geq\gamma^{h-2}$. The constant $\gamma>1$ will be chosen below to be large enough. Note that, in this way, the supports of $\chi_{0,i,+}$ and $\chi_{0,i,-}$ do not overlap. We define $\widehat{g}_{i,\omega}^{(\leq 0)}({\bf k})=\chi_{0,i,\omega}({\bf k})\widehat{g}_{i}({\bf k})$ and
@@
with $\widehat{g}^{(1)}({\bf k})=(1-\sum_{\omega}\chi_{0,i,\omega}(k))\widehat{g}_{i}({\bf k})$; this induces the Grassmann variable decomposition $\widehat{\psi}_{i,{\bf k},\alpha}=\widehat{\psi}_{i,{\bf k},\alpha}^{(1)}+\sum_{\omega=\pm}\widehat{\psi}^{(\leq 0)}_{i,{\bf k},\alpha,\omega}$ with propagators given by $\widehat{g}^{(1)}_{i}$ and $\widehat{g}^{(\leq 0)}_{i,\omega}$ respectively. Note that $\widehat{\psi}^{(1)}$ correspond to fermions with momenta far from the Fermi points, while $\widehat{\psi}^{(\leq 0)}$ with momenta around $K_{\pm}$.
@@
where $\widehat{g}_{i,\omega}^{(h)}({\bf k}):=f_{h,i,\omega}({\bf k})\widehat{g}_{i,\omega}^{(\leq 0)}$ in which $f_{h,i,\omega}:=\chi_{h,i,\omega}-\chi_{h-1,i,\omega}$ is a smooth cutoff function supported in $\gamma^{h-3}\leq|{\bf k}-{\cal K}_{\omega}^{i}|\leq\gamma^{h-1}$ such that $\sum_{h=-\infty}^{0}f_{h,i,\omega}=\chi_{0,i,\omega}$. The integration is done recursively in the following way: assume that we have integrated the fields $\psi^{(1)},..,\psi^{(h-1)}$ obtaining
$e^{W(\phi)}=\int\bar{P}(d\psi^{(\leq h)})e^{V^{(h)}(\psi,\phi)}$ (43)
@@
$\sum_{i,\omega,\omega^{\prime},m,\alpha,\alpha^{\prime}}\int_{\widehat{\Lambda}_{i}}d{\bf k}W^{(h,\omega,\omega^{\prime})}_{i,1,m,\alpha,\alpha^{\prime}}({\bf k})\psi_{i,{\bf k},\alpha,\omega}^{+}\psi_{1,{\bf k}+mb^{\prime},\alpha^{\prime},\omega^{\prime}}^{-}.$
@@
${\cal L}W^{(h,\omega,\omega^{\prime})}_{i,j,l}({\bf k})=\delta_{\omega,\omega^{\prime}}\delta_{i,j}\delta_{l,0}[W^{(h,\omega,\omega)}_{i,i,0}(0,K^{i}_{\omega})+$
$k_{0}\partial_{0}W^{(h,\omega,\omega)}_{i,i,0}(0,K^{i}_{\omega})+(k-K^{i}_{\omega})\partial W^{(h,\omega,\omega)}_{i,i,0}(0,K^{i}_{\omega}).$ (45)
@@
${\cal L}_{2}V^{(h)}:=\sum_{i,j,\omega,\alpha,\alpha^{\prime}}\int_{\widehat{\Lambda}_{i}}d{\bf k}z_{h,\omega,\alpha,\alpha^{\prime},i,j}(\mathbf{k}-\mathbf{K}^{i}_{\omega})_{j}\psi^{+}_{{\bf k},\omega,i,\alpha}\psi^{-}_{{\bf k},\omega,i,\alpha^{\prime}}$ (48)
@@
$\nu_{h,\omega,\alpha,\alpha^{\prime},i}:=\gamma^{-h}W_{i,i,0,\alpha,\alpha^{\prime}}^{h,\omega,\omega}(0,K^{i}_{\omega})$ (49)
$z_{h,\omega,\alpha,\alpha^{\prime},i,j}:=-\partial_{\mathbf{k}_{j}}W_{i,i,0,\alpha,\alpha^{\prime}}^{h,\omega,\omega}(0,K^{i}_{\omega}).$ (50)